# The Implication Problem for Functional Dependencies and Variants of Marginal Distribution Equivalences

### v1

@@ 5. Complexity of the implication problem for FDs+UMIs+UMDEs
This section considers the the implication problem for FDs+UMIs+UMDEs. We show that this implication problem also has a polynomial-time algorithm. Denote by $D$ the set of the variables appearing in $\Sigma\cup\{\sigma\}$, and suppose that $\Sigma$ is partitioned into the sets $\Sigma_{\text{FD}}$, $\Sigma_{\text{UMI}}$, and $\Sigma_{\text{UMDE}}$ that consist of the FDs, UMIs, and UMDEs from $\Sigma$, respectively.
First, note that if $\sigma$ is a UMI, it suffices to check whether $\sigma\in\mathsf{cl}(\Sigma_{\text{UMI}})$ as the inference rules for FDs and UMDEs do not produce new UMIs. Notice that $\Sigma_{\text{UMI}}$ can be viewed as an undirected graph $G=(D,\approx)$ such that each $x\approx y\in\Sigma_{\text{UMI}}$ means there is an undirected edge between $x$ and $y$. (We may assume that each vertex has a self-loop and all the edges are undirected, even though the corresponding UMIs might not be explicitly listed in $\Sigma_{\text{UMI}}$.) Suppose that $\sigma=x\approx y$. Then $\sigma\in\mathsf{cl}(\Sigma_{\text{UMI}})$ iff $y$ is reachable from $x$ in $G$. This is the problem of undirected s-t connectivity, which is known to be L-complete (Reingold, 2008), and thus it is in polynomial-time.
Hence, we can assume that $\sigma$ is an FD or a UMDE. The idea behind the following part of the algorithm is again based on the algorithm (Cosmadakis et al., 1990) for the implication problem for FDs and UINDs. In the FD+UMI case in Section 4, we could use the algorithm for FD+UIND and just slightly modify the set $\Sigma$, but since now we are also dealing with UMDEs, so we have to make some modifications.
The implication problem for FDs is known to be linear-time computable by the Beeri-Bernstein algorithm (Beeri and Bernstein, 1979). Given a set $\Delta\cup\{\mathop{=\!}(\bar{x},\bar{y})\}$ of FDs, the Beeri-Bernstein algorithm computes the set $\mathsf{fdclosure}(\bar{x},\Delta)=\{z\mid\Delta\models\mathop{=\!}(\bar{x},z)\}$. Then $\Delta\models\mathop{=\!}(\bar{x},\bar{y})$ holds iff $\mathsf{var}(\bar{y})\subseteq\mathsf{fdclosure}(\bar{x},\Delta)$. The Beeri-Bernstein algorithm keeps two lists: $\mathsf{fdlist}(\bar{x})$ and $\mathsf{attrlist}(\bar{x})$. The set $\mathsf{fdlist}(\bar{x})$ consists of FDs and is updated in order to keep the algorithm in linear-time (for more details, see (Beeri and Bernstein, 1979)). The set $\mathsf{attrlist}(\bar{x})$ lists the variables that are found to be functionally dependent on $\bar{x}$. The algorithm [^3] This presentation of the algorithm is based on (Cosmadakis et al., 1990). is as follows:
@@
Let $x\in D$, and define $\mathsf{desclist}(x)=\{y\in D\mid\Sigma\models\mathop{=\!}(x,y)\}$. We will construct an algorithm, called $\mathsf{Closure}(x,\Sigma)$, that computes the set $\mathsf{desclist}(x)$ by utilizing the Beeri-Bernstein algorithm. If we view the variables $D$ and the atoms from $\mathsf{cl}(\Sigma)$ as a multigraph of Definition 3.1, the set $\mathsf{desclist}(x)$ consists of the variables that are red descendants of $x$. Since FDs and UMIs interact only via UMDEs, there is no need to construct the whole multigraph. We only consider the UMIs in the initialization step, as they imply UMDEs. The algorithm $\mathsf{Closure}(x,\Sigma)$ keeps and updates $\mathsf{fdlist}(y)$ and $\mathsf{attrlist}(y)$ for each $y\in D$. When the algorithm halts, $\mathsf{attrlist}(x)=\mathsf{desclist}(x)$. The algorithm is as follows:
@@
In each iteration, before running the Beeri-Bernstein algorithm (step (4)), we find the strongly connected components (step (1)) and check all pairs of vertices for red edges and add red and blue edges if needed (step (2)). The time required for finding the strongly connected components is linear in the number of vertices and edges (Tarjan, 1972), so it takes $O(|D|+|\Sigma_{\text{FD}}|+n^{2}+|\Sigma_{\text{UMI}}|+|\Sigma_{\text{UMDE}}|)$, i.e. $O(n^{2})$ time. Going through the pairs of vertices in step (2) also takes $O(n^{2})$ time. Thus each iteration takes $O(n^{3}+n|\Sigma_{\text{FD}}|)$ time in total.
An iteration proceeds to step (4) only if new red edges are added in step (2). In each iteration, if new red edges are added, they merge together two or more red cliques in the same strongly connected component. In the beginning we have at most $n$ such cliques, so the algorithm cannot take more than $O(n)$ iterations. The total running time of $\mathsf{Closure}(x,\Sigma)$ is then $O(n^{4}+n^{2}|\Sigma_{\text{FD}}|)$.
If $\sigma=\mathop{=\!}(x,y)$, then $\Sigma\models\sigma$ iff $y\in\mathsf{desclist}(x)$. If $\sigma$ is a non-unary FD, in order to check whether $\Sigma\models\sigma$, it suffices to check whether $\Sigma_{\text{FD}}\cup\Delta\models\sigma$, where $\Delta=\{\mathop{=\!}(y,\bar{z})\mid y\in D,\mathsf{var}(\bar{z})=\mathsf{desclist}(y)\}$. This, again, can be done using the Beeri-Bernstein algorithm. Since the order of the variables $\mathsf{var}(\bar{z})$ in the tuple $\bar{z}$ does not matter for each FD $\mathop{=\!}(y,\bar{z})$, we may assume that the set $\Delta$ contains only one FD for each $y\in D$. Thus this step only adds $O(|\Sigma_{\text{FD}}|+n)$ extra time. (Note that constructing the set $\Delta$ does not affect the time-bound because $\mathsf{Closure}(x,\Sigma)$ actually computes $\mathsf{desclist}(y)$ for all $y\in D$.)
If $\sigma=x\approx^{*}y$, we can use the blue (undirected) subgraph of $G$ to check whether $\Sigma\models\sigma$. After the algorithm $\mathsf{Closure}(x,\Sigma)$ has halted, we only need to check whether $y$ is reachable from $x$ in the subgraph, which is an L-complete problem, as noted before.
Therefore, the above algorithm proves the following theorem:
Theorem 5.1 ().
@@ 6. Conclusion
We have presented a sound and complete infinite axiomatization for functional dependence, unary marginal identity, and unary marginal distribution equivalence, and shown that there is no finite axiomatization. The class of these dependencies was also shown to have Armstrong relations because in the completeness proof, we constructed an Armstrong relation for any set of FDs, UMIs, and UMDEs. We also showed that FD+UMI implication can be simulated with FD+UIND implication, and therefore it is in polynomial time. The implication problem for the full class of these dependencies was also shown to be in polynomial-time by constructing an algorithm. The following questions remain open:
@@
As for the second question, the implication problems for probabilistic independence and marginal identity have already been studied separately: the problem for probabilistic independence has been axiomatized in (Geiger et al., 1991), and the problem for marginal identity (over finite probability distributions) has recently been axiomatized in (Hannula and Virtema, 2021). Obtaining an axiomatization for both of these dependencies together is appealing, given how commonly IID assumption is used in probability theory and statistics. For probabilistic independence and unary marginal identity and distribution equivalence, the interaction seems simple, but as in the as in the case of functional dependence, adding nonunary marginal identities and distribution equivalences complicates things.
In the third question, we have both functional dependence and probabilistic independence together with marginal identity and marginal distribution equivalence. In the qualitative case, the decidability of the implication problem for independence atoms and (nonunary) functional dependencies seems to be open, but the axiomatization for independence atoms, unary functional dependencies and unary inclusion dependencies was introduced in (Hannula and Link, 2018). This suggests that it might be useful to consider just unary functional dependencies with probabilistic independence and unary marginal identity and distribution equivalence. This already makes the interaction more interesting than in the case of only probabilistic independence and unary marginal identity and distribution equivalence because we would need the cycle rules.
@@ Acknowledgements.
The author was supported by the Finnish Academy of Science and Letters (the Vilho, Yrjö and Kalle Väisälä Foundation) and by grant 345634 of the Academy of Finland. I would like to thank the anonymous referees for valuable comments regarding the previous version of the paper, and Miika Hannula and Juha Kontinen for useful discussions and advice.

### v2

@@ 5. Complexity of the implication problem for FDs+UMIs+UMDEs
In this section, we consider the implication problem for FDs+UMIs+UMDEs. We show that this implication problem also has a polynomial-time algorithm. Denote by $D$ the set of the variables appearing in $\Sigma\cup\{\sigma\}$, and suppose that $\Sigma$ is partitioned into the sets $\Sigma_{\text{FD}}$, $\Sigma_{\text{UMI}}$, and $\Sigma_{\text{UMDE}}$ that consist of the FDs, UMIs, and UMDEs from $\Sigma$, respectively.
First, note that if $\sigma$ is a UMI, it suffices to check whether $\sigma\in\mathsf{cl}(\Sigma_{\text{UMI}})$ as the inference rules for FDs and UMDEs do not produce new UMIs. Notice that $\Sigma_{\text{UMI}}$ can be viewed as an undirected graph $G=(D,\approx)$ such that each $x\approx y\in\Sigma_{\text{UMI}}$ means there is an undirected edge between $x$ and $y$. (We may assume that each vertex has a self-loop and all the edges are undirected, even though the corresponding UMIs might not be explicitly listed in $\Sigma_{\text{UMI}}$.) Suppose that $\sigma=x\approx y$. Then $\sigma\in\mathsf{cl}(\Sigma_{\text{UMI}})$ iff $y$ is reachable from $x$ in $G$. This can be checked in linear-time by using a breadth-first search algorithm.
Hence, we now treat the case where $\sigma$ is an FD or a UMDE. The idea behind the following part of the algorithm is again based on the algorithm (Cosmadakis et al., 1990) for the implication problem for FDs and UINDs. In the FD+UMI case in Section 4, we could use the algorithm for FD+UIND and just slightly modify the set $\Sigma$, but since now we are also dealing with UMDEs, so we have to make some modifications.
The implication problem for FDs is known to be linear-time computable by the Beeri-Bernstein algorithm (Beeri and Bernstein, 1979). Given a set $\Delta\cup\{\mathop{=\!}(\bar{x},\bar{y})\}$ of FDs, the Beeri-Bernstein algorithm computes the set $\mathsf{fdclosure}(\bar{x},\Delta)=\{z\mid\Delta\models\mathop{=\!}(\bar{x},z)\}$. Then $\Delta\models\mathop{=\!}(\bar{x},\bar{y})$ holds iff $\mathsf{var}(\bar{y})\subseteq\mathsf{fdclosure}(\bar{x},\Delta)$. (Note that the tuple $\bar{x}$ can be empty.) The Beeri-Bernstein algorithm keeps two lists: $\mathsf{fdlist}(\bar{x})$ and $\mathsf{attrlist}(\bar{x})$. The set $\mathsf{fdlist}(\bar{x})$ consists of FDs and is updated in order to keep the algorithm in linear-time (for more details, see (Beeri and Bernstein, 1979)). The set $\mathsf{attrlist}(\bar{x})$ lists the variables that are found to be functionally dependent on $\bar{x}$. The algorithm [^2] This presentation of the algorithm is based on (Cosmadakis et al., 1990). is as follows:
@@
Note that the Beeri-Bernstein algorithm can be run incrementally, i.e., it can first examine a set of FDs and find their closure and then, if new FDs are added afterwards, it can continue computing the closure with respect to this larger set of FDs. (Cosmadakis et al., 1990)
Let $x\in D$, and define $\mathsf{desclist}(x)=\{y\in D\mid\Sigma\models\mathop{=\!}(x,y)\}$. We will construct an algorithm, called $\mathsf{Closure}(x,\Sigma)$, that computes the set $\mathsf{desclist}(x)$ by utilizing the Beeri-Bernstein algorithm. If we view the variables $D$ and the atoms from $\mathsf{cl}(\Sigma)$ as a multigraph of Definition 3.2, the set $\mathsf{desclist}(x)$ consists of the variables that are red descendants of $x$. Since FDs and UMIs interact only via UMDEs, it is sufficient to construct the multigraph only for FDs and UMDEs. We only consider the UMIs in the initialization step, as they imply UMDEs. The algorithm $\mathsf{Closure}(x,\Sigma)$ keeps and updates $\mathsf{fdlist}(y)$ and $\mathsf{attrlist}(y)$ for each $y\in D$. When the algorithm halts, $\mathsf{attrlist}(x)=\mathsf{desclist}(x)$. The algorithm is as follows:
@@
In each iteration, before running the Beeri-Bernstein algorithm (step (4)), we find the strongly connected components (step (1)) and check all pairs of vertices for red edges and add red and blue edges if needed (step (2)). The time required for finding the strongly connected components is linear in the number of vertices and edges (Tarjan, 1972), so it takes at most $O(|D|+n^{2}+|\Sigma_{\text{UMI}}|+|\Sigma_{\text{UMDE}}|)$, i.e. $O(n^{2})$ time. Going through the pairs of vertices in step (2) also takes $O(n^{2})$ time.
An iteration proceeds to step (4) only if new red edges are added in step (2). In each iteration, if new red edges are added, they merge together two or more red cliques in the same strongly connected component. In the beginning we have at most $n$ such cliques, so the algorithm cannot take more than $O(n)$ iterations. Since the Beeri-Bernstein algorithm can operate incrementally, the total running time of $\mathsf{Closure}(x,\Sigma)$ is then $O(n^{3}+n|\Sigma_{\text{FD}}|)$.
If $\sigma=\mathop{=\!}(x,y)$, then $\Sigma\models\sigma$ iff $y\in\mathsf{desclist}(x)$. If $\sigma$ is a constant atom or a non-unary FD, in order to check whether $\Sigma\models\sigma$, it suffices to check whether $\Sigma_{\text{FD}}\cup\Delta\models\sigma$, where $\Delta=\{\mathop{=\!}(y,\bar{z})\mid y\in D,\mathsf{var}(\bar{z})=\mathsf{desclist}(y)\}$. This, again, can be done using the Beeri-Bernstein algorithm. Since the order of the variables $\mathsf{var}(\bar{z})$ in the tuple $\bar{z}$ does not matter for each FD $\mathop{=\!}(y,\bar{z})$, we may assume that the set $\Delta$ contains only one FD for each $y\in D$. Thus this step only adds $O(|\Sigma_{\text{FD}}|+n)$ extra time. (Note that constructing the set $\Delta$ does not affect the time-bound because $\mathsf{Closure}(x,\Sigma)$ actually computes $\mathsf{desclist}(y)$ for all $y\in D$.)
If $\sigma=x\approx^{*}y$, we can use the blue (undirected) subgraph of $G$ to check whether $\Sigma\models\sigma$. After the algorithm $\mathsf{Closure}(x,\Sigma)$ has halted, we only need to check whether $y$ is reachable from $x$ in the subgraph, which is a linear-time problem, as noted before.
Therefore, in the worst case the above algorithm takes $O(n^{3}+n|\Sigma_{FD}|)$ time and we have the following theorem:
Theorem 5.1.
@@
Note that the time-bound of the algorithm is the same as that of the algorithm for the implication problem for FDs+UINDs in (Cosmadakis et al., 1990).
@@ 6. Conclusion
We have presented a sound and complete infinite axiomatization for functional dependence, unary marginal identity, and unary marginal distribution equivalence, and shown that there is no finite axiomatization. The class of these dependencies was also shown to have Armstrong relations. We also showed that FD+UMI implication can be simulated with FD+UIND implication, and therefore it is in polynomial time. The implication problem for the full class of these dependencies was also shown to be in polynomial-time by constructing an algorithm. The following questions remain open:
@@
As for the second question, the implication problems for probabilistic independence and marginal identity have already been studied separately: the problem for probabilistic independence has been axiomatized in (Geiger et al., 1991), and the problem for marginal identity (over finite probability distributions) has recently been axiomatized in (Hannula and Virtema, 2021). Obtaining an axiomatization for both of these dependencies together is appealing, given how commonly IID assumption is used in probability theory and statistics. For probabilistic independence and unary marginal identity and distribution equivalence, the interaction seems simple, but as in the case of functional dependence, adding nonunary marginal identities and distribution equivalences introduces additional complexity.
In the third question, we have both functional dependence and probabilistic independence together with marginal identity and marginal distribution equivalence. In the qualitative case, the decidability of the implication problem for independence atoms and (nonunary) functional dependencies is still open, but the axiomatization for independence atoms, unary functional dependencies and unary inclusion dependencies was introduced in (Hannula and Link, 2018). This suggests that it might be useful to consider just unary functional dependencies with probabilistic independence and unary marginal identity and distribution equivalence. This already makes the interaction more interesting than in the case of only probabilistic independence and unary marginal identity and distribution equivalence because we would need the cycle rules.
@@ Acknowledgements.
The author was supported by the Finnish Academy of Science and Letters (the Vilho, Yrjö and Kalle Väisälä Foundation) and by grant 345634 of the Academy of Finland. I would like to thank all the anonymous referees for valuable comments regarding both this and the previous conference version of the paper, and Miika Hannula and Juha Kontinen for useful discussions and advice.